OpenStudy (anonymous):

lara chooses a square number she rounds it up to the nearest hundred her answer is 200

OpenStudy (anonymous):

So we know our answer is in between \[\sqrt{200} and \sqrt{100}\] but not equal to it because she had to round. So we know our x whatever it is, is greater than 10 then right?

OpenStudy (anonymous):

And so we know to find the first number that when you square it, is greater than 200. What is that?

OpenStudy (anonymous):

\[14^{2}=196\] so I'm guessing 14 or \[\sqrt{196}\]

OpenStudy (anonymous):

You got it! 14. You square it and rounded up the nearest hundred is 200. 13 also works for this interestingly enough since its nearest hundred is also 200 but I would go with 14. nice job!